# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

### v2

@@ 2.1 Depth: The Hardest Problem Sampled in RLVR
where is the binary reward of rollout, is the mean value of the group rewards and is the standard deviation of the group rewards . In the case of binary rewards, also represents the accuracy of LLM rollouts. Dr. GRPO (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. We show more results of std-based advantage in Appendix B.1.
@@ 3.1 Difficulty Adaptive Rollout Sampling (DARS)
As illustated in Figure 6, given a data batch of reasoning questions, DARS operates in two phases: (i) pre-rollout difficulty estimation step that assigns to each question a scalar difficulty score ; and (ii) multi-stage rollout re-balancing step that dynamically decides how many additional trajectories shall be allocated to so that the cumulative advantage for low-accuracy problems is up-weighted. In order to simplify the subsequent formula representation, we define
Phase 1: Pre-Rollout Difficulty Estimation.
For every , we draw a light first-stage rollout consisting of independent trajectories . Let the per-trajectory reward be binary, . We define the empirical accuracy
@@
Phase 2: Multi-Stage Rollout Re-Balancing.
Let denote the cumulative advantage under GRPO for a group whose average accuracy is with rollout size . We aim to reallocate additional trajectories across the mini-batch so that the effective cumulative advantage for each question becomes an increasing function of its difficulty. To control the computational cost of rollout sampling, we set the maximum sampling upper limit . To this end, we design two re-balancing schedules.
@@ 4.1 Setup
We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23. We combine all of the evaluation benchmarks to report Pass@1 (Avg@128) and Pass@K performance. The training data used in this work is OpenR1-45K, which is a subset of OpenR1-Math-220k (Hugging Face, 2025). More details are shown in Appendix A.1.
@@ 4.3 Training Dynamics and Ablation Study
Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1. Figure 8 plots Pass@128 against Pass@1 under a variety of experimental settings. It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings. For Pass@32-Pass@1 training dynmaics, please refer to Appendix B.2
@@ 6 Conclusion
Appendix A Training and Evaluation Details
A.1 Setting Details
@@
Appendix B More Experimental Results
B.1 Ablation Study on std-based Advantage Computation
@@
B.2 Depth and Breadth Synergy for Pass@1 and Pass@32
@@
B.3 Comparison of ET/HW Schedule in Breadth Scaling
@@
Appendix C Discussion and Future Work
@@
C.1 Hyperparameter Control of Cumulative Advantage Shape
@@
C.2 Potential Pass@K to Pass@1 Training Transition

### v3

@@ 2.1 Depth: The Hardest Problem Sampled in RLVR
where is the binary reward of rollout, is the mean value of the group rewards and is the standard deviation of the group rewards . In the case of binary rewards, also represents the accuracy of LLM rollouts. Dr. GRPO (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. We show more results of std-based advantage in Appendix C.1.
@@ 3.1 Difficulty Adaptive Rollout Sampling (DARS)
As shown in Figure 6, given a data batch of reasoning questions, DARS operates in two phases: (i) pre-rollout difficulty estimation that assigns to each question a scalar difficulty score ; and (ii) multi-stage rollout re-balancing that dynamically decides how many additional trajectories shall be allocated to so that the cumulative advantage for low-accuracy problems is up-weighted. To simplify the subsequent formula representation, we define
Phase 1: Pre-Rollout Difficulty Estimation. For every , we draw a light first-stage rollout consisting of independent trajectories . Let the per-trajectory reward be binary, . We define the empirical accuracy
@@
Phase 2: Multi-Stage Rollout Re-Balancing. Let denote the cumulative advantage under GRPO for a group whose average accuracy is with rollout size . We aim to reallocate additional trajectories across the mini-batch so that the effective cumulative advantage for each question becomes an increasing function of its difficulty. To control the computing cost, we cap the rollout sampling upper limit at . To this end, we design two rebalancing schedules.
@@ 4.1 Setup
We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23. We combine all of the evaluation benchmarks to report Pass@1 (Avg@128) and Pass@K performance. The training data used in this work is OpenR1-45K, which is a subset of OpenR1-Math-220k (Hugging Face, 2025). More details are shown in Appendix B.1.
@@ 4.3 Training Dynamics and Ablation Study
Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1. Figure 8 plots Pass@128 against Pass@1 under a variety of experimental settings. It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings.
@@ 6 Conclusion
Appendix A Derivation of Additional Rollouts
The cumulative advantage for a group with accuracy and total rollout size is given by:
where .
After the first-stage rollout of size , the initial cumulative advantage is:
Our goal is to determine the number of additional trajectories needed so that the final cumulative advantage meets a target value .
Equal-Treatment (ET) Schedule:
The target cumulative advantage is set to be constant for all questions with :
We solve for :
The rollout size must be an integer, and we cap the total rollout sampling upper limit at , so
Hardness-Weighted (HW) Schedule:
The target cumulative advantage increases with difficulty:
We solve for :
Again, using the baseline advantage notation , we obtain:
Appendix B Training and Evaluation Details
B.1 Setting Details
@@
Appendix C More Experimental Results
C.1 Ablation Study on std-based Advantage Computation
@@
C.2 Depth and Breadth Synergy for Pass@1 and Pass@32
@@
C.3 Comparison of ET/HW Schedule in Breadth Scaling
@@
Appendix D Discussion and Future Work
@@
D.1 Hyperparameter Control of Cumulative Advantage Shape
@@
D.2 Potential Pass@K to Pass@1 Training Transition